# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v2

@@ Abstract
How can cognitive science build generalizable theories that span the full scope of natural situations and behaviors? We argue that progress in Artificial Intelligence (AI) offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes. We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms, and models that accommodate them, may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize. First, we review cases from cognitive science and neuroscience where naturalistic paradigms elicit distinct behaviors or engage different processes. We then discuss recent progress in AI that shows that learning from naturalistic data yields qualitatively different patterns of behavior and generalization, and discuss how these findings impact the conclusions we draw from cognitive modeling, and can help yield new hypotheses for the roots of cognitive and neural phenomena. We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding. We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition—together with a reductive understanding of the processes and principles by which they do so.
Short Abstract
How should cognitive science pursue generalizable theories? We argue that progress in Artificial Intelligence (AI) offers opportunities to embrace naturalistic experiments and computational models that can accommodate them. We first review literature suggesting that building generalizable theories may require a broader range of naturalistic experimental paradigms. We then suggest that progress in AI enables incorporating naturalistic paradigms and models without losing experimental control or theoretically-grounded understanding. We argue that we are entering a new era in which cognitive science can build computational models that solve the real problems of natural cognition, and derive a reductive understanding of those solutions.
Keywords: Artificial intelligence, benchmarks, cognitive science, engineering, generalizability, naturalistic, task-performing models, theory
@@ 1 Introduction
Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition  (McClelland,, 2009). Cognitive models can range from simplified models of a high-level process (e.g. Frank and Goodman,, 2012; Cohen et al.,, 1990), to task-performing models of a particular domain (Newell,, 1973), or beyond (Newell,, 1994). Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001; Tenenbaum et al.,, 2011), cognitive architectures (e.g. Ritter et al.,, 2019; Laird,, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles. However, despite progress, there is still substantial debate on the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g. Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Nastase et al.,, 2020; Wise et al.,, 2023; Cisek and Green,, 2024). Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
@@
The path we propose is motivated by recent practical and conceptual developments in AI. In the last ten years, AI has progressed substantially in building models that can perform a wide range of tasks in various domains. We now have vision models that can classify and segment real-world images without external supervision (He et al.,, 2020; Caron et al.,, 2021), reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and language models that can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022). Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al.,, 2014; Schrimpf et al.,, 2021). These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ; Kanwisher et al.,, 2023), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023). Yet they have also led to debate over how we test our models and interpret the results (Bowers et al.,, 2023; DiCarlo et al.,, 2023; Dentella et al.,, 2024). Thus, it remains an open question if and how the progress in AI can contribute to addressing the challenges of cognitive science.
@@
Figure 1: Naturalistic computational cognitive science: the what, why, and the how. The first section (§2) provides an overview of “naturalistic computational cognitive science”. In (§34), we motivate using naturalistic experimental paradigms and learning-based approached for cognitive research. The remaining sections (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories. (Figures reproduced from Saxe et al.,, 2019; Geirhos et al.,, 2018; Doshi and Konkle,, 2023.)
@@
In this work, we attempt to weave a thread of arguments that synthesizes these literatures, spanning from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility. We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve a deeper understanding of the full range of natural intelligence. We present an overview of this framework and our arguments in Figure 1.
@@
Our paper proceeds as follows. First, in §2, we unpack what we mean by “naturalistic” computational cognitive science, by providing a pragmatic outline of the features that we believe deserve more emphasis in our experimental paradigms and models. We argue that researchers should expand their task paradigms to more closely approximate the breadth of the settings in which their theoretical constructs should be expected to generalize—first, by adding relevant variables that may interact with the theoretical constructs in question, and second by incorporating more variability within existing parameters. Likewise, researchers should expand their models to more faithfully approximate the computational task that the natural system is solving.
@@
In the next two sections, we motivate these goals by arguing that increasing naturalism is necessary for developing generalizable understanding. First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can yield different behavior, and alter how tasks engage computational mechanisms or neural systems. For example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes. Other work has found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020; Hasson et al.,, 2015) and more are changed when naturalistic elements are introduced. These works suggest that experiments that span a broader range of naturalistic settings can be essential for drawing correct inferences about the system as a whole. Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can yield qualitatively different patterns of generalization than models trained in simplified settings. These differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena. Together, these two sections of our paper argue that pursuing naturalistic experiments, and models that accommodate them, may be necessary for achieving a complete understanding of cognition.
In the remainder of the paper, we turn to the practice of naturalistic computational cognitive science, and its contributions to building generalizable cognitive theories. First, in §5, we outline key lessons from AI in developing generalizable task-performing models for increasingly naturalistic data conditions: the benefits of being able to frictionlessly reproducing research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions. Finally, in §6, we outline how complicated experiments and opaque models can contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021). We close in §7 by discussing the broader context and implications of naturalistic computational cognitive science.
1.1 What this article is not
We first address some potential misconceptions we anticipate about our arguments.
First, we do not intend to argue that simple experimental paradigms are useless. Rather, we argue that incorporating more naturalistic settings alongside simpler ones is essential to developing more generalizable cognitive theories.
We also do not intend to promote a particular class of models, aside from our emphasis on learning. We believe that our arguments are compatible with many learning-based cognitive models, e.g. Bayesian cognitive models (Lee and Wagenmakers,, 2013), reinforcement learning (Collins,, 2019), predictive representations (Carvalho et al.,, 2024), program synthesis models (Rule et al.,, 2024), etc.
Finally, we do not advocate for black-box foundation models trained directly to reproduce experimental data as the goal of naturalistic computational cognitive science. While this can be a useful tool (e.g. for experimental design), we advocate for theory-driven development of predictive models that capture human behavior as an important tool in the quest to develop and validate generalizable cognitive theories.
In summary, this paper is not intended to reject the current experimental, modeling, and theoretical approaches of cognitive science. Rather, we advocate for extending from simpler experiments to more naturalistic ones. We then advocate for building task-performing models that can predict human behavior across these paradigms, and linking these models to simpler reductive models and theories of that behavior. We argue that this approach will enable building more generalizable theories that explicitly link cognitive capacities to the full scope of natural behavior to which those capacities contribute.
Figure 2: Overview of naturalistic computational cognitive science. Given a novel computational model, researchers generate predictions to test their model in a simplified setting. If this model successfully predicts human behavior, researchers see if their model continues to generalize human predictions in increasingly naturalisic conditions.
@@ 2 What is “naturalistic” computational cognitive science?
Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict and understand human behavior across increasingly naturalistic settings. This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions. In machine learning (§5), analogous strategies have enabled models that can work with real-world data, albeit imperfectly.
As in other approaches within computational cognitive science, this approach starts with formulating models and designing controlled experiments to isolate core phenomena. Like prior work, we place an emphasis on model generalization, i.e. predicting human behavior on data the model hasn’t been exposed to—including data from new experimental conditions or tasks (Busemeyer and Wang,, 2000). Therefore,
@@
One key facet of naturalistic computational cognitive science is that we place emphasis on having models—in particular, the same model—predict human behavior across increasingly naturalistic conditions that cover a broader space of settings in which the theoretical construct would be expected to generalize. We see two main paths toward increasing naturalism in experimental paradigms:
@@
We provides examples in Figure 3.
We describe a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize. We describe a model as more naturalistic if it is capable of generalizing its predictions across a broader range of contexts. Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments (see §6.1).
@@
  • Large state spaces that reflect the complexity of real-world environments (Wise et al.,, 2023).
  • Incorporating other social agents that learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
  • A continually changing environment (Abel et al.,, 2024).
@@
  • Architectures that operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g. low-dimensional or discrete inputs containing only task-relevant features).
  • Architectures with naturalistic action spaces—for example, modeling motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
@@
  • Unsupervised (Higgins et al.,, 2021), self-supervised (Konkle and Alvarez,, 2022), or intrinsic (Chentanez et al.,, 2004; Oudeyer et al.,, 2013) learning algorithms.
  • Social (Henrich,, 2016) or cultural (Cook et al.,, 2024) learning algorithms.
@@
  • Prospective algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
In summary, naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of the environment, and (2) develop cognitive theories that unite these models with reductive understanding. In the coming sections, we will first motivate how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to how these goals can be achieved.
3 The benefits of naturalistic experimental settings
Figure 4: Overview of the benefits of increasingly naturalisic experimental conditions.
In this section, we outline the benefits of naturalistic experimental settings for cognitive science. Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby alter the scientific inferences that we draw. We review examples where naturalistic experimental settings have changed our understanding of a system, or have helped to disentangle competing models that simpler settings could not.
@@ 3 The benefits of naturalistic experimental settings
As an orienting conceptual example, imagine that we are doctors studying heart function, and we chose to only study it while the participants are lying down and resting. We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study processes like how respiration affects heart rate. However, this control and reliability comes precisely from removing the many factors of natural variation that affect the heart, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc. Studying heart function across a broader range of naturalistic settings would be necessary to understand the system in full, and how the physiological (and psychological) processes involved interact. For example, if we experimentally manipulated breathing across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine its effect on heart function, and how that effect varies across contexts. By studying phenomena across varied naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.
3.1 Behavior can be different in naturalistic experiments
Many cognitive analyses begin by studying a task designed to isolate some cognitive capability. However, in many cases task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted setting unless we also consider a broader range of variation. In this section, we review some examples illustrating changes in behavior or capabilities in naturalistic settings.
Example 1: moral behavior can change when moving from textual vignettes to virtual reality. Recent research shows that behavioral choices can change when moving from text to virtual reality, by driving a shift from moral judgment (i.e. deciding whether things are moral) to moral action (Francis et al.,, 2017, see Figure 4 A). When participants were presented the classic trolley dilemma in text form, only 40% said they would pull a switch to divert the trolley, thus killing one person to save five others. However, when the same scenario was presented in virtual reality, where participants must simulate the action, 55% chose to pull the switch. The difference was even more pronounced in the “footbridge” variant of the dilemma. In the text version, only 10% of participants said they would push someone off a bridge to save five others. However, in virtual reality, 65% of participants choose to push the person. These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.

### v4

@@ Abstract
How can cognitive science build generalizable theories that span the full scope of natural situations and behaviors? We argue that progress in Artificial Intelligence (AI) offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes. We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms, and models that accommodate them, may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize. We review cases from cognitive science and neuroscience where naturalistic paradigms elicit distinct behaviors or engage different processes. We then discuss recent progress in AI that shows that learning from naturalistic data yields qualitatively different patterns of behavior and generalization, and examine how these findings impact the conclusions we draw from cognitive modeling, and can help yield new hypotheses for the roots of cognitive and neural phenomena. We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding. We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition, together with a reductive understanding of the processes and principles by which they do so.
Keywords: Artificial intelligence, benchmarks, cognitive science, ecological validity, engineering, generalizability, naturalistic, task-performing models, theory
@@ 1 Introduction
Cognitive scientists build models to make our theories concrete—which offers testable predictions, eliminates ambiguities (Guest and Martin, 2021), and can reveal unexpected properties of cognition (McClelland, 2009). Cognitive models can range from simplified models of a high-level process (e.g. Frank and Goodman, 2012; Cohen et al., 1990), to task-performing models of a particular domain (Newell, 1973), or beyond (Newell, 1994). Many cognitive paradigms—connectionism (McClelland et al., 1986), Bayesian inference (Tenenbaum and Griffiths, 2001; Tenenbaum et al., 2011), cognitive architectures (e.g. Ritter et al., 2019; Laird, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles. However, despite progress, there is still substantial debate on the generalizability of our theories (Yarkoni, 2022; Eckstein et al., 2021), whether our models and theories are adequately constrained (e.g. Jones and Love, 2011; Rahnev and Denison, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Nastase et al., 2020; Wise et al., 2023; Cisek and Green, 2024). Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
@@
The path we propose is inspired by recent practical and conceptual developments in AI. We now have vision models that can classify and segment real-world images without external supervision (He et al., 2020; Caron et al., 2021), reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Adaptive Agent Team et al., 2023), and language models that can solve many language-specified tasks (Radford et al., 2019; Wei et al., 2022). Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al., 2014; Schrimpf et al., 2021). This has prompted debate on whether such complex models can be explanatory (Cao and Yamins, 2024a; Kanwisher et al., 2023), what they imply for our theories (Hasson et al., 2020; Perconti and Plebe, 2020; Piantadosi, 2023), and a broader question of how we should test and interpret our theories generally (Bowers et al., 2023; DiCarlo et al., 2023; Dentella et al., 2024). Thus, whether and how this progress can serve cognitive science remains open.
@@
Figure 1: Naturalistic computational cognitive science: the what, why, and the how. The first section (§2) provides an overview of “naturalistic computational cognitive science”. In (§34), we motivate using naturalistic experimental paradigms and learning-based approaches for cognitive research. The remaining sections (§56) focus on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories. (Figures reproduced from Saxe et al., 2019; Geirhos et al., 2018; Doshi and Konkle, 2023.)
@@
In this work, we attempt to weave a thread of arguments that synthesizes these literatures, ranging from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility. We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve a deeper understanding of the full range of natural intelligence. We present an overview of the paper and our arguments in Figure 1.
@@
Our paper proceeds as follows. In §2, we specify what naturalistic computational cognitive science means in practice. We argue that researchers should broaden their paradigms beyond the settings in which their theoretical constructs are typically tested—by both adding variables that interact with those constructs and by sampling more broadly within existing parameters. Likewise, we should build models that more faithfully approximate the computational task that the natural system is solving.
@@
The next two sections address why. We argue that increasing naturalism is necessary—not merely beneficial—for developing a generalizable theory of cognition. In §3, we review evidence from neuroscience, cognitive science, and AI showing that naturalistic paradigms yield qualitatively different behavior and engage different computational mechanisms—processes spanning visual processing (Amme et al., 2024), learning (Rosenberg et al., 2021; Collins, 2024), memory (Helbing et al., 2020; Hasson et al., 2015), and moral reasoning (Francis et al., 2017) all shift when naturalistic elements are introduced. These findings suggest that naturalistic experiments are essential for building a complete theory of the system. In §4, we go further: training on naturalistic data does not merely reveal different mechanisms but can also shape them, producing qualitatively different generalization patterns than training in simplified settings. These differences can alter the inferences we draw from our models and shed new light on the origins of cognitive and neural phenomena. Together, these two sections establish that pursuing naturalistic experiments—and models that accommodate them—is necessary for a complete understanding of cognition (Figure 2).
The remaining sections address the how. In §5, we distill a key lesson from AI for developing task-performing models in naturalistic settings: building generalizable models requires iterative and collective effort. We propose dynamic meta-benchmarks—living collections that researchers continually expand—as the coordination mechanism, and outline a development strategy for fast, frictionless iteration. In §6, we show how naturalistic experiments and complex models can anchor generalizable cognitive theories—uniting task-performing predictive models with reductive understanding of underlying mechanisms, using tools such as rational analysis (Anderson, 1990) and new approaches to interpreting complex models (Geiger et al., 2021). We close in §7 with the broader context and implications of naturalistic computational cognitive science.
@@ 2 What is “naturalistic” computational cognitive science?
Figure 2: Two examples of naturalistic computational cognitive science. Given a computational theory, researchers generate initial predictions to test their theory in a simplified setting. Theories are validated by their ability to predict human behavior in increasingly naturalistic conditions, and refined where they fail to predict correctly. (Top) Progression of theories of sequential task generalization. (Bottom) Examples of progression in understanding PRC involvement in difficult visual judgments.. Figures adapted from Momennejad et al. (2017); Tomov et al. (2021); Carvalho et al. (2025); Bonnen et al. (2021, 2026).
Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict and understand human behavior across increasingly naturalistic settings. This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions. Of course, our emphasis on naturalism does not imply that simple experimental paradigms are useless—they remain valuable, and the goal is to extend them with naturalistic settings rather than replace them. In machine learning (§5), analogous strategies have enabled models that can work with real-world data, albeit imperfectly.
Aside from our focus on learning-based methods, this approach is model-class agnostic and compatible with many learning-based frameworks, including Bayesian models (Lee and Wagenmakers, 2013), RL (Collins, 2019), predictive representations (Carvalho et al., 2024), and program synthesis (Rule et al., 2024). The main criterion we advocate for is that models are theory-driven—not black-box foundation models trained to reproduce experimental data (Binz et al., 2024), though these may be a useful tool for experimental design.
As in other approaches within computational cognitive science, this approach starts with formulating models and designing controlled experiments to isolate core phenomena. Like prior work, we place an emphasis on model generalization, i.e. predicting human behavior on data the model hasn’t been exposed to—including data from new experimental conditions or tasks (Busemeyer and Wang, 2000). Therefore,
@@
One key facet of naturalistic computational cognitive science is that we place emphasis on having models—in particular, the same model—predict human behavior across increasingly naturalistic conditions that cover the space of settings in which the theoretical construct is expected to generalize. We see two main paths toward increasing naturalism in experimental paradigms (and provide examples in Figure 3):
@@
We describe a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize. We describe a model as more naturalistic if it is capable of generalizing its predictions across a broader range of contexts. Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments testing a hypothesis (see §6.1).
@@
  • Large state spaces that reflect the complexity of real-world environments (Wise et al., 2023).
  • Incorporating other social agents that learn with the agent (Hu et al., 1998) and that the organism must interact with (Vélez and Gweon, 2021).
  • A continually changing environment (Abel et al., 2024).
@@
  • Architectures that operate over sensory inputs like natural images (Yamins et al., 2014), speech (Kell et al., 2018), or natural language (Schrimpf et al., 2021) rather than simplified stimuli (e.g. low-dimensional or discrete inputs containing only task-relevant features).
  • Architectures with naturalistic action spaces—for example, modeling motor control and embodiment rather than abstract, symbolic actions (Merel et al., 2019).
@@
  • Unsupervised (Higgins et al., 2021), self-supervised (Konkle and Alvarez, 2022), or intrinsic (Chentanez et al., 2004; Oudeyer et al., 2013) learning algorithms.
  • Social (Henrich, 2016) or cultural (Cook et al., 2024) learning algorithms.
@@
  • Prospective algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al., 2013; De Silva et al., 2023).
These four dimensions are not a definition of naturalism. They are examples of directions along which a theory of cognition must hold up to count as general. Generalizability means extending along axes like these—and often theories are not extended along any.
3 The importance of naturalistic experimental settings for building generalizable theories
Figure 4: Overview of the benefits of increasingly naturalistic experimental conditions.
@@ 3 The importance of naturalistic experimental settings for building generalizable theories
As an orienting conceptual example, let us imagine that we are doctors studying heart function, and have chosen to study participants lying at rest. The setting offers high test-retest reliability and a controlled view of processes like how respiration affects heart rate. But that control comes precisely from removing the variation the heart evolved to handle — rapid adaptation to movement, stress, and the like. Understanding the system in full, and how its physiological and psychological processes interact, requires studying the heart across more naturalistic settings. For instance, manipulating breathing across activities like sitting, speaking, and running would pin down its effect on heart function more sharply than rest alone, and reveal how that effect varies with context. Varied naturalistic settings are how we elaborate the system and build complete theories of its function.
Here, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can help to disentangle underlying mechanisms, and achieve more generalizable theoretical understanding.
Example 1: mechanisms for credit assignment are only exposed when actions have delayed consequences. In naturalistic settings, people learn how actions lead to consequences hundreds of timesteps away. Consider navigating a university campus: you learn that heading north leads to the music school, then discover a jazz cafe just off that path. Credit assignment (Sutton and Barto, 2018) is what lets you recognize the northward route serves both. Prior work modeled this with successor features (Gershman, 2018; Tomov et al., 2021), but only studied human behavior in worlds of around 13 states—small enough that models can succeed without any explicit long-horizon mechanism. In larger worlds, successor features predict fast reaction times but cannot handle nearby-object transfer. In contrast, model-based planning handles this transfer but predicts slow deliberation. Humans display both. To explain this, Carvalho et al. (2025) proposed multitask preplay: people mentally rehearse potential detours in advance and cache the results for fast retrieval later on (Figure 2, top).
Testing multitask preplay required a large environment: in smaller worlds, planning is cheap enough that cached retrieval looks no different from deliberation. Implementing cached retrieval with a brain-plausible learning rule then made the credit-assignment problem concrete—the agent must credit a northward choice (taken to reach the music school) for the cafe discovered along the way. Such cross-task transfer is a known failure mode of temporal-difference learning: approximation errors inflate the value of actions that should not be taken. Multitask preplay required additional machinery to mitigate this, pointing to mechanisms the brain may use for the same job. Because the model was developed in a naturalistic setting, it continued to predict human behavior in an even more naturalistic one—a large, partially observable 2D world with many objects.
Example 2: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models. A long-running debate asks whether perceptual “oddity” tasks (identifying the object that differs from the others in a set) recruit the Medial Temporal Lobe (MTL). Murray et al. (2007) and others argue that MTL is recruited to discriminate “complex” stimulus sets that visual cortex cannot resolve alone, so MTL lesions selectively impair performance on those stimuli; others find no such interaction and attribute it to methodological artifacts (Buffalo et al., 1998; Suzuki, 2009). Bonnen et al. (2021); Bonnen and Eldridge (2023) resolved the impasse by passing all the raw stimuli from prior experiments through pretrained convolutional vision models and testing whether the oddity could be picked out in representation space. When the vision model alone could distinguish the oddity, MTL was dispensable; when it could not, MTL was required. Some “complex” stimulus sets from prior experiments fell on each side of this split. “Complexity” had been underspecified: differences in operationalization that conceptual descriptions abstracted away drove the seemingly-conflicting findings. This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms—resolving otherwise intractable theoretical disagreements.
Example 3: Different learning mechanisms only yield distinct predictions under working memory load. One challenge of overly-simplified tasks is potential aliasing of different solutions, where many different computational approaches yield the same behavior. Collins (2024) presents one example where a behavior ostensibly implemented as a standard reinforcement learning algorithm may instead be produced from simpler value-free associative learning combined with working memory. [^1] Whether this alternative counts as an alternative implementation of RL is orthogonal to our argument; our point is that the models yield distinct predictions in only some experimental regimes. Collins notes that “Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the set of stimuli within learning blocks to a handful of objects, rather than only two or three—i.e., expanding the range of set sizes more clearly disentangles the learning processes. Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them. By contrast, by imposing a broad range of evaluations on the system (cf. Nau et al., 2024)—especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence. More constraints on the mapping make it less ambiguous, as is highlighted in the contravariance principle of Cao and Yamins (2024b).
3.2 Behavior can be qualitatively different in naturalistic experiments
Cognitive analyses typically isolate a capacity with a task designed to elicit it. But behavior often shifts with seemingly-orthogonal aspects of that task, and theories tuned to one setting risk overfitting to it. Three cases illustrate the risk concretely—putatively well-isolated capacities (memory, learning, moral choice) behave qualitatively differently under naturalistic demands.